OpenStudy (anonymous):

Find the solution to the system of equations by using either graphing or substitution. y=6-x and y=x-2 a.(2,4) b.(-4,2) c.(4,2) d.no solutions

OpenStudy (anonymous):

y = 6 - x now sub 6 - x in for y in the other equation y = x - 2 6 - x = x - 2 -- subtract x from both sides -x - x + 6 = -2 -- subtract 6 from both sides -x - x = -2 - 6 -- combine like terms -2x = -8 -- divide both sides by -2 x = -8/-2 x = 8/2 when reduced = 4 now sub 4 in for x in either equation to find y y = x - 2 y = 4 - 2 y = 2 check... y = 6 - x 2 = 6 - 4 2 = 2 (correct) x = 4 and y = 2

OpenStudy (jim766):

the other thing you could do since the both equal y is make them equal to each other y = 6-x and y = x-2 6-x = x-2 since they both equal y they must = each other 2x = 8 x = 4 then sub back in the 4 to find y

OpenStudy (jim766):

you get the same answer, just another way to do it